OpenStudy (anonymous):

Jamie has a deck of 60 sports cards, of which some are baseball cards and some are football cards. Jamie pulls out a card randomly from the deck, records its type, and replaces it in the deck. Jamie has already recorded six baseball cards and nine football cards. Based on these data, what is, most likely, the number of baseball cards in the deck?

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

sry this looks higher level probablity than me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@niels5x9 can you help

OpenStudy (mathmath333):

6x+9y=60 where x = bascket ball cards and y = football cards

OpenStudy (mathmath333):

so i think it simplifies to \(2x+3y=20\) we have to check this number for x and y \(\begin{align}\{x,y\}=\{1,1\}\\ \{2,2\}\\ \{3,3\}\\ \{4,4\}\end{align} \)

OpenStudy (anonymous):

substitute x and y?

OpenStudy (mathmath333):

we see that \( \{4,4\}\) satisfies hence \(4\times 6=24\) basket ball cards, your answer is correct

OpenStudy (anonymous):

@mathmath333 thanks